OpenStudy (anonymous):

Identify the 25th term of the arithmetic sequence 2, 1 3/5, 1 1/5...

OpenStudy (alexwee123):

an = a1 + (n – 1)d

OpenStudy (campbell_st):

the 1st term a = 2, common difference is -2/5 n = 25 use the formula \[T _{n}=a + (n -1)d\] substitute and evaluate to find the answer